OpenStudy (anonymous):

which is the slowest rate ? a)2,133km in 6hr b)3,267 km in 11r c)1456km in 4hr d)2461 km in 9 hr

OpenStudy (anonymous):

C because if you 456/4=114 a 2133/6=355.5 b 3267/11=297 d 2461/9=273.4

OpenStudy (goformit100):

c)1456km in 4hr is correct.

OpenStudy (anonymous):

Great